OpenStudy (anonymous):

What is the x-intercept of the line with the equation -2x+1/2y=18 ? ( , 0)

OpenStudy (cwrw238):

the x intercept is found by putting y = 0 and solving for x

OpenStudy (anonymous):

okay so how do you do that lol

OpenStudy (cwrw238):

so we have -2x + 1/2(0) = 18

OpenStudy (cwrw238):

and 1/2 (0) = ?

OpenStudy (cwrw238):

I have replaced y in the equation by 0

OpenStudy (anonymous):

\[\-2x+\cancel{\frac{1}{2}y}=18\\ -2x=18\]

OpenStudy (anonymous):

i got -9

OpenStudy (cwrw238):

thats right

OpenStudy (cwrw238):

the intercept is (-9,0)

OpenStudy (anonymous):

oh okay thank you to both of you

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

could you help me with one more

OpenStudy (anonymous):

@cwrw238 or @satellite73

OpenStudy (cwrw238):

sorry gotta go right now

OpenStudy (anonymous):

ok thats fine